OpenStudy (anonymous):

6 – 4x = 4(3 – x) – 6 Help!! Me!!

OpenStudy (anonymous):

distribute the 4 to each term inside the ( ). Let me know what you get. a (x + n) = ax + an

OpenStudy (anonymous):

no solution?

OpenStudy (anonymous):

6-4x+4x-12=0 6-12=0 -6

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

ur wlc

OpenStudy (anonymous):

No -6 cannot equal 0

OpenStudy (anonymous):

so that persons wrong? @ChmE

OpenStudy (anonymous):

6 – 4x = 4(3 – x) – 6 6 - 4x = 12 - 4x - 6 6 - 4x = 6 - 4x x = x infinite amount of solution

OpenStudy (anonymous):

Yes, I posted the correct above

OpenStudy (anonymous):

okay thanks alot

OpenStudy (anonymous):

x can be any number. Infinite amount of solutions

OpenStudy (anonymous):

@Muskan

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

okay i get it now